OpenStudy (anonymous):

if f(x)=x^1/2-x and g(x)=2x^3-x^1/2, find f(x)-g(x)

OpenStudy (anonymous):

i would really appreciate it if someone would actually help me this time!...

OpenStudy (anonymous):

ok ill help u

OpenStudy (anonymous):

thank you somuch!

OpenStudy (anonymous):

this would be your equation (x^1/2-x)-(2x^3-x^12)

OpenStudy (anonymous):

Answers are: -2x^3-2x-2x^1/2 -2x^3-2x 2x^3-2x+2x^1/2 -2x^3+2x^1/2

OpenStudy (anonymous):

okay... my is like my worst subject so, could you explain how to do that? please

OpenStudy (anonymous):

math*

OpenStudy (anonymous):

is this what the problem looks like?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

??...

OpenStudy (anonymous):

umm hmm this one is tough

OpenStudy (anonymous):

yeah ik DX

OpenStudy (firekat97):

you know that we can write x^1/2 as √x right?

OpenStudy (anonymous):

i didnt know that

OpenStudy (firekat97):

well you should try and remember that its pretty useful :)

OpenStudy (firekat97):

so if we have f(x) - g(x) we can write this as (√x - x) - (2x^3 - √x)

OpenStudy (firekat97):

do you know how you would simplify - (2x^3 - √x)?

OpenStudy (anonymous):

thank you so much, math is my worst subject but im almost done with the class. thanks for helping!! (:

OpenStudy (anonymous):

let me try, but most likely no

OpenStudy (anonymous):

nope i dont

OpenStudy (firekat97):

okay so imagine there is a 1 after the - sign so you have to simplify -1(2x^3 - √x) do you know the FOIL method?

OpenStudy (anonymous):

i forgot it havent done it in a while

OpenStudy (anonymous):

is it -2x^3+2x^1/2

OpenStudy (anonymous):

?

OpenStudy (firekat97):

yup! thats correct!

OpenStudy (anonymous):

omg thanks so much!!

OpenStudy (anonymous):

would you mind helping me out i have three more questions left?

OpenStudy (firekat97):

so now we have x^1/2 - x - 2x^3 + x^1/2 and we want to simplify this by collecting like terms

OpenStudy (firekat97):

were nearlyyyy finished with this question ahahha

OpenStudy (anonymous):

oh haha i thought we were alredy done lol i already went on to the next question

OpenStudy (firekat97):

lol dw we're close

OpenStudy (anonymous):

can we skip that one ?

OpenStudy (anonymous):

A machine is set up to cut metal strips of varying lengths and widths based on the time (t) in minutes. The change in length is given by the function l(t)=t^2 - sq. root of t, and the change in width is given by w(t)=t^2 - 2t 1/2. Which function gives the change in area of the metal strips?

OpenStudy (firekat97):

okay do you know formula for the area of a rectangle?

OpenStudy (anonymous):

a=wl

OpenStudy (anonymous):

Answers: a(t)=t^4+2t a(t)=t^4+2t+3t^5/2 a(t)=t^4-3t^5/2+2t a(t)=t^4+2t-2t^1/2+square root of t

OpenStudy (firekat97):

yup so we want to do l(t) x w(t) do you see why?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

wait yes i do